OpenStudy (anonymous):

x + y = 4 x - 2y = -5 If the first equation is multiplied by 2 and then the equations are added, the result is

OpenStudy (anonymous):

3x= 13 3x= 3 x= 3 possible answers^

OpenStudy (saifoo.khan):

x = 1 y =3

OpenStudy (anonymous):

Which answer?

OpenStudy (saifoo.khan):

both are the answers.

OpenStudy (anonymous):

3x= 13? 3x= 3? x= 3?

OpenStudy (saifoo.khan):

whn u divide 3 by 3 the answer is 1.

OpenStudy (anonymous):

so x= 3?

OpenStudy (saifoo.khan):

x = 1

OpenStudy (anonymous):

I need to know which answer i pick.

OpenStudy (sciencyguy):

Take it step by step...if you try to rush right to the answer it's easier to make mistakes. x + y = 4 so multiplying that by two gets you to the new equation 2x + 2y = 8. Add that to x - 2y = -5. Now you have 2x + x + 2y - 2y = 8 - 5 That's 3x = 3; divide both sides by 3 and you get x = 1... Then you can put your answer for x back into either equation and get the answer for y. Take it slow!

OpenStudy (saifoo.khan):

x = 1 look, whn u divide 3 by 3, the answer is 1 so x = 1